OpenStudy (anonymous):

please help me, i am struggling with math the sum of two numbers is 7. one number is three more than the other. find the smaller number. a.-4 b.2 c.4 d.8

OpenStudy (anonymous):

Hi there! Do you know how to solve systems of linear equations?

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

are you willing to?

OpenStudy (anonymous):

Let two numbers be x and y x + y = 7 x = y + 3 x = 7 - x + 3 2x = 10 x = 5

OpenStudy (anonymous):

OK. We can probably do this without it, actually. let's call 'x' the smaller number. What's the value of the larger number?

OpenStudy (anonymous):

can you show me how you got the answer? please?

OpenStudy (anonymous):

b.2

OpenStudy (anonymous):

x + x + 3 = 7 2x + 3 = 7 2x = 4 x = 2

OpenStudy (anonymous):

x=5 x+y = y y = 2 Answer is b.

OpenStudy (anonymous):

One number is 3 larger than the other, right? So if 'x' is the smaller number, how can we write the larger number, in terms of x?

OpenStudy (anonymous):

ktklown - you are a good person!

OpenStudy (anonymous):

haha, thank you :-)

OpenStudy (anonymous):

You are trying to teach. I was trying to do same. Not always what is needed. Giving answer is what is needed.

OpenStudy (anonymous):

For people who are just trying to get the answer I try to send them to wolfram alpha so we can spend time here actually trying to teach

OpenStudy (anonymous):

But it is pointless just to give out answer

OpenStudy (anonymous):

x+y=7 x+(3+x)=7 2x+3=7 x = (7 -3)/2 x= 2

OpenStudy (sasogeek):

GT giving the answer is stupidity, useless and baseless if the person doesn't understand

OpenStudy (anonymous):

I understand. Lot of times you are ignored if you try to teach.

OpenStudy (anonymous):

but sometimes they are in a test or quiz and they are in a hurry

OpenStudy (anonymous):

The people who ignore efforts to teach, I tell to go to wolfram, or I ignore, or if they're being mean I give them the wrong answer

OpenStudy (anonymous):

If they are in a hurry, they should fail.

OpenStudy (anonymous):

"they're being mean I give them the wrong answer" don't waste time doing that

OpenStudy (sasogeek):

I don't help ppl to fail and neither would i help them to cheat, if what i tell u turns out to be wrong, you'll blame me... u need to know why i'm writing what i did... that's how you learn. don't give out answers if the person is in a hurry. lets say the person is in the middle of a test... giving the person an answer because the person is in a hurry means helping the person to cheat.... you wouldn't give the answer if u were there with the person in the exam room...